a proton moving with a speed of 10^6m/s collides with a second proton initial at rest.one proton emerge from the collision at the angle 30 and other 60 degree from the incident direction.calculate the speed of proton.
1
Expert's answer
2020-04-08T10:26:04-0400

Let us assume the mass of proton as 'm'

As collision takes place so momentum will be conserved here

"p_i=p_f\\p_i=m\times10^6+m\times0"

Let us see the final case one proton is at 30o and other one is at 60o

in y direction final momentum will be zero

"mv_1\sin30^o=mv_2\sin60^o\\v_1=v_2\sqrt3" ..............................(1)


now let's see the momentum in the x direction

"mv_1\cos30^o+mv_2\cos60^o=m\times10^6\\"

substituting the value of "v_1" in the above equation

"v_2\times\frac{3}2+v_2\times\frac12=10^6\\v_2=5\times10^5m/s"


from eq. (1)

we get "v_1=5\sqrt3\times10^5m/s"